# COLT: Enhancing Video Large Language Models with Continual Tool Usage

### v2

@@ Introduction
We additionally compare the proposed COLT with popular continual learning methods on VideoToolBench test split to demonstrate the life-long tool-usage learning capability of our method.
@@
Evaluation Metrics. To evaluate how the system retains tool-using knowledge over time, we adapt the conventional continual learning metric (Wang et al. 2024a; Chaudhry et al. 2018) to our tool-using scenario. We set up two metrics: 1) average accuracy of a specific tool denotes the overall call accuracy during the incremental learning process; 2) average forgetting is defined as the mean difference between the maximum tool call accuracy throughout the past learning process and the current tool call accuracy. A reduction in the value of average forgetting is indicative of enhanced continual learning capability. Refer to supplementary materials for the formula and detailed explanations.
@@
Compared Methods. We set comparison experiments as follows: 1) Sequential training: training on new data with the pre-trained weights on previous tool data as initialization; 2) Rehearsal training: replay past tool data (i.e., “buffer”) and combine them with new data. Buffer size is respectively set to 10/30/50 for each tool in experiments; 3) Popular continual learning methods including L2P (Wang et al. 2022c), Dual (Wang et al. 2022b) and CODA (Smith et al. 2023).
Performance Analysis. We conduct experiments in two settings, i.e., five/ten groups with two/one tools per group. The comparison results in Table 6 reveal that our proposed COLT consistently exhibits superior performance in both average accuracy and forgetting across both settings. For example, COLT outperforms CODA (Smith et al. 2023) by 3.5% on average accuracy under the group number of five, highlighting its adeptness in mitigating the deleterious catastrophic forgetting.
Ablation Studies
@@
We conduct extensive ablation studies to provide more insights into our proposed COLT. The experiments are conducted on MVBench (Li et al. 2023c) with the continual learning model variant COLT $_{\text{5$\times$2}}$.
@@
Ablation on Training Strategies. Our COLT adopts a three-stage training pipeline. We conduct ablation studies on the training process by respectively skipping the second and third stages of training. The comparison results on MVBench (Li et al. 2023c) in Table 4(a) demonstrate that both stages are crucial to the final performance, e.g., skipping the second stage leads to a 1.4% drop in average scores.
@@
Ablation on Training Loss. We use the straight-through estimator (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) for the training of stage 2 and stage 3 (c.f. Eq. (5)), which contains the quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$ to enables the mutual updates between the selected tool prompts and query features. The ablation results in Table 4(b) underscore the significance of both quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$.
@@
Ablation on Prompt Selection. Recall that we select the most matched tool prompts based on the similarity between the textual instruction features and codebook (i.e., text-codebook). Here we ablate on the prompt selection based on the vision-codebook similarity or the average similarities of both. The results are listed in Table 4(c), which showcases that the text-codebook similarity is more reliable. That could be attributed to that it is easier for video LLMs to decide whether to invoke or which specific tool to invoke from the user instructions instead of the input visual data.
@@
Ablation on Tool Prompt Position. We ablate on the insertion position of the tool prompt. We’ve listed three options including tool-vision-text, vision-tool-text, and vision-text-tool, demonstrating placing the tool feature before/between/after the vision and text features, respectively. The findings in Table 4(d) indicate that the performance is insensitive with regard to the tool prompt position.
Table 7: Ablations on hyper-parameters including the codebook size NN and the number of selected prompt $K$.
@@
Figure 4: Accuracy of zero-shot video-question answering on MSVD vs. continually introduced tasks. The specific tool name for each abbreviation is available in Table 1.
Ablation on hyper-parameters. We conduct ablation studies on the codebook size NN and the number of selected prompt $K$. The comparison results are listed in Table 7. As shown, the average performance is positively correlated with the codebook size NN and reaches saturation at $T=50$. The optimal results are achieved when setting $K=3$.
Visualizations of performance vs. continually introduced tasks. To intuitively show the impact of incrementally introducing tools, we show the phased zero-shot video-question answering performance under different training strategies. Specifically, we report the accuracy score on MSVD-QA dataset in Figure 4. As shown, equipping video LLMs with dense video caption tools leads to the significant performance boost for all four training strategies. Besides, our proposed COLT shows better learning stability, i.e., less forgetfulness when faced with new tool data.
Conclusions
In this work, we present COLT which continually learns new tool-using knowledge in a data-stream scenario for general-purpose video understanding. To achieve this, we propose a learnable tool codebook where the specific tool is retrieved and activated according to the similarities with the user instructions. Due to the absence of a video-centric tool-using instruction-tuning dataset within the community, we devised VideoToolBench to address this deficiency and foster the exploration of tool-using capacities of video LLMs. Experimental results indicate that our proposed COLT adeptly invokes the necessary tools with precision, thereby achieving state-of-the-art performance on widely used benchmarks and the proposed VideoToolBench.
Appendix A1 Appendix
@@ Conclusions
  • Dataset and Evaluation Metric Details
    • Dataset Structure of VideoToolBench
    • Prompt for Curating VideoToolBench
@@
    • Quality Control of VideoToolBench
    • Evaluation Metrics of Tool Continual Learning
@@
Dataset and Evaluation Metric Details
Dataset Structure of VideoToolBench. Each instance in VideoToolBench contains two rounds of conversation between human and gpt. The response of gpt includes three fields of thought, action and value. One example is available in Table A1.
Table A1: Dataset structure of VideoToolBench.
An example of VideoToolBench dataset
@@
Sure thing! I’$l$ initiate a object segmentation model to fulfill your request. Please wait while the processing takes place.
@@
Prompt for Curating VideoToolBench. We collect VideoToolBench by self-instruct GPT-3.5-turbo. The prompt and in-context learning cases are shown in Table A2.
Table A2: System prompts and in-context learning examples for the generation of VideoToolBench.
@@
You are an AI assistant and you receive a set of conversations in json string format, the content of which is used as data for instruction finetune. Your task is to rephrase the dialogue to generate a new example. Note to give it in standard json format. DO NOT modify the “from” and “actions” parts. Only modify the “value” and “thoughts” part.
@@
{”from”: ”human”, ”value”: ”Transcribe the audio content from the provided video.”, ”actions”: []}, {”from”: ”gpt”, ”thoughts”: ”I need to find an ASR model and initiate the transcription process as per the user’s request.”, ”actions”: [{”API_\_name”: ”asr”, ”API_\_params”: {}}], ”value”: ”Certainly! I’$l$ employ an Automatic Speech Recognition (ASR) model. Please wait while the ASR model processes the video”}, {”from”: ”human”, ”value”: ”asr output: You have used an ASR model to transcribe the audio content from the provided video based on my original request.”, ”actions”: []}, {”from”: ”gpt”, ”thoughts”: ”Now that the ASR model has processed the video, I can update the user on the transcription results.”, ”actions”: [], ”value”: ”The ASR transcription you requested has been completed. Here is the transcribed text for your reference.”}
@@
Instruction Tuning Dataset Reformat. In the second and third stage of COLT training, we reformat the instruction tuning dataset from LLaVA v1.5 (Liu et al. 2023a) and Video-ChatGPT (Maaz et al. 2023) into the thought-action-value pattern for the data format conformity. Specifically, the thought is generated by GPT indicating that the question can be answered without invoking any tools: The questions can be answered by the information in the context, without need any external tools. The action is an empty list while value is the original response.
@@
Quality Control of VideoToolBench: To ensure the reliability and integrity of the proposed dataset, we conduct quality control from two primary aspects: data format checks and of semantic meaning checks. The format check entails verifying whether the generated data adheres to a predefined structure, specifically confirming that each entry corresponds to a two-round conversation and that the responses from the GPT model contain essential components such as thought/action/value keys. Furthermore, this phase involves confirming the consistency of tool names used throughout the dataset. Manual checks on semantic meanings are conducted to validate the contextual relevance and accuracy of the generated content, ensuring that it aligns with the intended purpose of the dataset and maintains coherence within the conversations.
Evaluation Metrics of Tool Continual Learning. We detail the metrics of average accuracy and average forgetting used in Sec. Comparisons with Continual Learning Methods.
$\text{AA}_{k}=\frac{1}{k}\sum_{j=1}^{k}a_{k,j}.$ (A1)
$f_{j}^{k}=\max_{l\in\{1,\cdots,k-1\}}a_{l,j}-a_{k,j},\quad\forall j<k.$ (A2)
$\text{AF}_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}f_{j}^{k}.$ (A3)
Experimental Results
In Sec. Training, we employ an additional CLIP-initialized encoder ${f}_{\boldsymbol{\beta}}$ to extract query feature $\mathbf{H}_{\mathbf{q}}$. Then in Eq. (2), the appropriate tool prompt is selected based on the cosine similarity with $\mathbf{H}_{\mathbf{q}}$. Here we conduct the ablation study which directly uses the encoded LLM features $\mathbf{H}_{\mathbf{w}}$ for similarity computation. The comparison results in Table A3 demonstrate that the introduced CLIP feature $\mathbf{H}_{\mathbf{q}}$ leads to better performance. This might be due to the fact that LLM features are more compatible with auto-regressive language modeling, whereas CLIP features are better suited for retrieval-oriented applications.
@@
Table A3: Ablations of the query feature source for tool prompt selection.
@@
Broader Impact
Qualitative Results
We present additional qualitative results on MVBench (Li et al. 2023c), zero-shot video-question answering (Chen and Dolan 2011; Xu et al. 2016; Yu et al. 2019), and VideoToolBench test split in Figure A1, Figure A2, and Figure A3, respectively. Notably, our COLT precisely captures dynamic video information and generate more reasonable responses.
Figure A1: Qualitative results on MVBench.
Figure A2: Qualitative results on zero-shot video-question answering.
Figure A3: Qualitative results on the test split of VideoToolBench.

### v3

@@ 5.3 Comparisons with Continual Learning Methods
We additionally compare the proposed COLT with popular continual learning methods on VideoTool test split to demonstrate the life-long tool-usage learning capability of our method.
@@
Evaluation Metrics. To evaluate how the system retains tool-using knowledge over time, we adapt the conventional continual learning metric wang2024comprehensive; chaudhry2018riemannian to our tool-using scenario. We set up two metrics: 1) average accuracy of a specific tool denotes the overall call accuracy during the incremental learning process; 2) average forgetting is defined as the mean difference between the maximum tool call accuracy throughout the past learning process and the current tool call accuracy. A reduction in the value of average forgetting is indicative of enhanced continual learning capability. Refer to the supplementary materials for the formula and detailed explanations.
@@
Compared Methods. We set comparison experiments as follows: 1) Sequential training: training on new data with the pre-trained weights on previous tool data as initialization; 2) Rehearsal training: replay past tool data (i.e., “buffer”) and combine them with new data. Buffer size is respectively set to 10/30/50 for each tool in experiments; 3) Popular continual learning methods including L2P wang2022learning, Dual wang2022dualprompt, and CODA smith2023coda.
Performance Analysis. We conduct experiments in two settings, i.e., five/ten groups with two/one tools per group. As shown in Table 5(b), COLT consistently exhibits superior performance in both average accuracy and forgetting across both settings. For example, under the five-group setting, COLT outperforms CODA smith2023coda by 3.5% in average accuracy, highlighting its effectiveness in mitigating catastrophic forgetting. Notably, despite being given rehearsal buffers of size 10/30/50 per tool, rehearsal-based methods remain significantly weaker (Table 5(b)), as they depend on storing and replaying past data and still struggle with interference across heterogeneous tools. In contrast, COLT achieves stronger retention without accessing old samples, demonstrating that its codebook-based tool memory provides a more stable and efficient continual-learning mechanism than rehearsal.
5.4 Ablation Studies
@@
We conduct extensive ablation studies to provide more insights into our proposed COLT. The experiments are conducted on MVBench li2023mvbench with the continual learning model variant COLT $_{\text{5$\times$2}}$.
@@
Ablation on Training Strategies. Our COLT adopts a three-stage training pipeline. We conduct ablation studies on the training process by respectively skipping the second and third stages of training. The comparison results on MVBench li2023mvbench in Table 4(a) demonstrate that both stages are crucial to the final performance, e.g., skipping the second stage leads to a 1.4% drop in average scores.
@@
Ablation on Training Loss. We use the straight-through estimator van2017neural; bengio2013estimating for the training of stage 2 and stage 3 (c.f. Eq. equation 5), which contains the quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$ to enables the mutual updates between the selected tool prompts and query features. The ablation results in Table 4(b) underscore the significance of both quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$.
@@
Ablation on Prompt Selection. Recall that we select the most matched tool prompts based on the similarity between the textual instruction features and codebook (i.e., text-codebook). Here we ablate on the prompt selection based on the vision-codebook similarity or the average similarities of both. The results are listed in Table 4(c), which showcases that the text-codebook similarity is more reliable. That could be attributed to the fact it is easier for video LLMs to decide whether to invoke or which specific tool to invoke from the user instructions instead of the input visual data.
@@
Ablation on Tool Prompt Position. We ablate on the insertion position of the tool prompt. We’ve listed three options, including tool-vision-text, vision-tool-text, and vision-text-tool, demonstrating placing the tool feature before/between/after the vision and text features, respectively. The findings in Table 4(d) indicate that the performance is insensitive with regard to the tool prompt position.
i Accuracy of zero-shot video-question answering on MSVD vs. continually introduced tasks. The specific tool name for each abbreviation is available in Table 1.
@@
$i$ Ablations on hyper-parameters including the codebook size $N$ and the number of selected prompt $K$.
Ablation on hyper-parameters. We conduct ablation studies on the codebook size $N$ and the number of selected prompts $K$. The comparison results are listed in Figure 4ii. As shown, the average performance is positively correlated with the codebook size $N$ and reaches saturation at $T=50$. The optimal results are achieved when setting $K=3$.
Visualizations of performance vs. continually introduced tasks. To intuitively show the impact of incrementally introducing tools, we show the phased zero-shot video-question answering performance under different training strategies. Specifically, we report the accuracy score on MSVD-QA dataset in Figure 4i. As shown, equipping video LLMs with dense video caption tools leads to the significant performance boost for all four training strategies. Besides, our proposed COLT shows better learning stability, i.e., less forgetfulness when faced with new tool data.
6 Limitations
While COLT demonstrates strong continual tool-use capabilities, several limitations remain. First, VideoTool is primarily GPT-generated, which may introduce distribution biases and limit the dataset’s real-world diversity compared with fully human-curated corpora. Second, the current dataset focuses mainly on single-tool and simple multi-tool compositions, which do not fully capture the complexity, interdependence, and noise present in real-world tool ecosystems, where tool chains may be substantially longer and more intricate. Future work includes extending VideoTool to more diverse and human-in-the-loop scenarios, scaling COLT to larger and more complex tool vocabularies, and exploring reinforcement learning or other adaptive strategies to support dynamic and automatic tool composition.
7 Conclusions
In this work, we present COLT, which continually learns new tool-using knowledge in a data-stream scenario for general-purpose video understanding. To achieve this, we propose a learnable tool codebook where the specific tool is retrieved and activated according to the similarities with the user instructions. Due to the absence of a video-centric tool-using instruction-tuning dataset within the community, we devised VideoTool to address this deficiency and foster the exploration of tool-using capacities of video LLMs. Experimental results indicate that our proposed COLT adeptly invokes the necessary tools with precision, thereby achieving state-of-the-art performance on widely used benchmarks and the proposed VideoTool test split.
Broader Impact Statement
Appendix A Appendix
@@
  • Dataset Details
    • Dataset Structure of VideoTool
    • Prompt for Curating VideoTool
@@
    • Quality Control of VideoTool
  • Illustrations of Tool Selection
@@
A.1 Dataset Details
Dataset Structure of VideoTool. Each instance in VideoTool contains two rounds of conversation between human and gpt. The response of gpt includes three fields of thought, action and value. One example is available in Table 6.
Table 6: Dataset structure of VideoTool.
An example of VideoTool dataset
@@ A.1 Dataset Details
Sure thing! I’ll initiate a object segmentation model to fulfill your request. Please wait while the processing takes place.
@@
Prompt for Curating VideoTool. We collect VideoTool by self-instruct GPT-3.5-turbo. The prompt and in-context learning cases are shown in Table 7.
Table 7: System prompts and in-context learning examples for the generation of VideoTool.
@@
You are an AI assistant and you receive a set of conversations in json string format, the content of which is used as data for instruction finetune. Your task is to rephrase the dialogue to generate a new example. Note to give it in standard json format. DO NOT modify the “from" and “actions" parts. Only modify the “value" and “thoughts" part.
@@
{"from": "human", "value": "Transcribe the audio content from the provided video.", "actions": []}, {"from": "gpt", "thoughts": "I need to find an ASR model and initiate the transcription process as per the user’s request.", "actions": [{"API_\_name": "asr", "API_\_params": {}}], "value": "Certainly! I’ll employ an Automatic Speech Recognition (ASR) model. Please wait while the ASR model processes the video"}, {"from": "human", "value": "asr output: You have used an ASR model to transcribe the audio content from the provided video based on my original request.", "actions": []}, {"from": "gpt", "thoughts": "Now that the ASR model has processed the video, I can update the user on the transcription results.", "actions": [], "value": "The ASR transcription you requested has been completed. Here is the transcribed text for your reference."}
@@
Instruction Tuning Dataset Reformat. In the second and third stage of COLT training, we reformat the instruction tuning dataset from LLaVA v1.5 (liu2023improved) and Video-ChatGPT (maaz2023video) into the thought-action-value pattern for the data format conformity. Specifically, the thought is generated by GPT indicating that the question can be answered without invoking any tools: The questions can be answered by the information in the context, without need any external tools. The action is an empty list while value is the original response.
@@
Quality Control of VideoTool: To ensure the reliability and integrity of the proposed dataset, we conduct quality control from two primary aspects: data format checks and of semantic meaning checks. The format check entails verifying whether the generated data adheres to a predefined structure, specifically confirming that each entry corresponds to a two-round conversation and that the responses from the GPT model contain essential components such as thought/action/value keys. Furthermore, this phase involves confirming the consistency of tool names used throughout the dataset. Manual checks on semantic meanings are conducted to validate the contextual relevance and accuracy of the generated content, ensuring that it aligns with the intended purpose of the dataset and maintains coherence within the conversations.
A.2 Illustrations of Tool Selection
Figure 5 shows the full pipeline of our tool selection mechanism.
Figure 5: Illustration of the tool selection mechanism based on cosine similarity. Given the query embedding $\mathbf{h}_{q}$ and a tool codebook $\{\mathbf{p}_{1},\dots,\mathbf{p}_{N}\}$, we first compute the cosine similarities between $\mathbf{h}_{q}$ and each tool prompt $\mathbf{p}_{i}$. The most relevant tools are then selected (e.g., via top-$K$) and concatenated with the visual/text embeddings before being fed into the large language model.
A.3 Experimental Results
In Sec. 4.2, we employ an additional CLIP-initialized encoder ${f}_{\bm{\beta}}$ to extract query feature $\mathbf{H}_{\mathbf{q}}$. Then in Eq. equation 2, the appropriate tool prompt is selected based on the cosine similarity with $\mathbf{H}_{\mathbf{q}}$. Here we conduct the ablation study which directly uses the encoded LLM features $\mathbf{H}_{\mathbf{w}}$ for similarity computation. The comparison results in Table 9 demonstrate that the introduced CLIP feature $\mathbf{H}_{\mathbf{q}}$ leads to better performance. In addition, we further analyze how tool-use training and the codebook mechanism affect general video understanding beyond tool-centric benchmarks. As reported in Table 8, removing VideoToolBench data results in performance close to standard video-VQA models, while discarding the codebook causes substantial degradation across all benchmarks. These results suggest that the codebook plays a critical role in preserving and reusing tool-related knowledge, enabling COLT to transfer tool-use behaviors to standard VQA tasks rather than merely memorizing tool traces.
Method MSVD-QA MSRVTT-QA ActivityNet-QA
Acc. $\uparrow$ Score $\uparrow$ Acc. $\uparrow$ Score $\uparrow$ Acc. $\uparrow$ Score $\uparrow$
COLT (w/o VideoToolBench) 70.9 3.8 58.9 3.4 45.4 3.2
COLT (w/o Codebook) 52.3 3.3 42.7 3.0 33.7 2.9
COLT-joint 78.2 4.2 65.1 3.6 54.7 3.8
Table 8: Controlled ablations on standard VQA benchmarks to disentangle the effects of tool-use training (VideoToolBench) and the codebook mechanism.
@@ A.3 Experimental Results
Table 9: Ablations of the query feature source for tool prompt selection.
@@
A.4 Qualitative Results
We present additional qualitative results on MVBench (li2023mvbench), zero-shot video-question answering (chen2011collecting; xu2016msr; yu2019activitynet), and VideoTool test split in Figure 6, Figure 7, and Figure 8, respectively. Notably, our COLT precisely captures dynamic video information and generate more reasonable responses.
Figure 6: Qualitative results on MVBench.
Figure 7: Qualitative results on zero-shot video-question answering.
Figure 8: Qualitative results on the test split of VideoTool.